sydneym:

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms?

surjithayer:

first term a=3 c.d. d-9-3=6 no. of terms =26 \[s _{n}=\frac{ n }{ 2 }[2a+(n-1)d]\] \[s _{26}=\frac{ 26 }{ 2 }[2(3)+(26-1)(6)]\] \[s _{26}=13[6+25\times 6]\] \[s _{26}=13\times 156\]\[=?\]

toga:

The first term of the arithmetic sequence is 3 and the common difference between the terms is 6. To find the sum of an arithmetic sequence, we can use the formula: S = n/2 [2a + (n-1)d] where S is the sum of the sequence, n is the number of terms, a is the first term, and d is the common difference. Plugging in the given values, we get: S = 26/2 [2(3) + (26-1)(6)] S = 13 [6 + 150] S = 13 x 156 S = 2028 Therefore, the sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms is 2028.

Manny300303199:

The sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms is 2028.